| Operation Summary for January 23, 2012 - January 25, 2012 |
|
Monday January 23 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, at 10:25 AM, the H- Source voltage dropped when its anode power supply tripped off; it reset. Operations established beam to the Meson FTBF experiment T-1018 at 10:26 AM. At 10:42 AM, a Linac quadrupole magnet power supply (L:QPS129) tripped off; it reset. At 12:06 PM, a Booster RF station (BRF4) tripped off; it reset. At 12:35 PM, a Switchyard (SY) person attempted to move a Meson collimator to a specific point, but the collimator moved all the way into the beam and got stuck; Instrumentation personnel investigated. At 12:48 PM, a Linac Klystron RF station (KRF5) tripped off; it reset. Operators prepared Meson enclosures M03 and 4 for access at 12:58 PM. MI personnel began a g-2 (g minus 2) study at 1:23 PM. Instrumentation personnel managed to get the Meson collimator out of the beam and Operations reestablished beam to T-1018 at 1:55 PM. A team of people made up from Operations and Water Group personnel began turning on pond water pumps at Tevatron sectors F1 and F2 at 2:46 PM to help stabilize F2 cooling for the SY LCW system; they also turned on pumps at sectors E4 and F3. During the evening shift, at 4:05 PM, a Booster vacuum pump (B:IPRF15) tripped off and brought down two RF stations (BRF3 and 4); everything recovered and reset. At 4:15 PM, L:QPS104 and 105 tripped off; they reset. The g-2 study ended at 4:43 PM. Operations disabled T-1018 beam at 10 PM. Operators reported that the H- Source was unstable for most of the shift. Tuesday January 24 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, Operators reported at 3:15 AM that the H- Source drop out and it took 30 minutes for the source to recover; Operators made adjustments. At 7 AM, Neutron Therapy Facility personnel began preparing to treat two patients later in the day. During the day shift, at 8:16 AM, MIRF5 tripped off and took down the NuMI beam permit; everything reset. The NuMI beam permit tripped off at 9:36 AM due to bad beam positions. An Instrumentation person reported at 10:04 AM that he was changing out the motor controller relays for the Meson MT4 collimator. KRF6 tripped off at 10:12 AM; it reset. At 10:30 AM, problems with the H- Source during NTF treatment required Operations to halt beam for about an hour. At 11:46 AM, Preaccelerator personnel accessed the H- Source to work on a gas valve, which halted all beam. During this access RF personnel worked on MIRF12, Booster personnel worked on their low-level RF phase lock electronics, a Linac person pulled cables through the Linac floor, a MI person investigated an ion pump (607B), and Fire technicians accessed the MI-60 power supply rooms to test equipment. At 12:02 PM, a NuMI person reported that the MI-62 LCW system had been warming up; Water Group personnel investigated. Operations established beam to MiniBooNE at 12:12 PM. Operations resumed sending beam to NuMI and T-1018 at 12:32 PM. At 12:34 PM, LRF3 tripped off; a Linac person made adjustments. At 1:32 PM, a Booster person halted beam for ten minutes to conduct some low-level RF work. At 1:43 PM, a Linac person halted beam for ten minutes to work on the H-Source. KRF3 tripped off at 2:57 PM; it reset. During the evening shift, Operators conducted a Linac tune at 5:56 PM. At 7:57 PM, B:IPRF15 tripped off and brought down BRF3 and 4; everything reset. Operations disabled T-1018 beam at 10:01 PM. Operators conducted another Linac tune at 11:55 PM. Wednesday January 25 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, at 1:49 AM, BRF11 tripped off; it reset. At 2:12 AM, the LRF5 vacuum pump tripped off; it recovered. LRF4 tripped off at 3:58 AM; it reset. At 5:38 AM, a MI kicker (KPS4) tripped off due to a Thyratron pre-fire (KPS4F); KPS4F would not reset. Operators consulted a MI person and then decided to run without the gap-clearing kickers until a kicker person could come in and investigate. He arrive at 6:43 AM and began his investigation. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with the plans of Booster and MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 01/25/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |